# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

### v4

@@ A.1 Question Answering
Question Answering (QA) [226] aims to give answers to the given queries, by generating multiple-choice or free-text responses. A multiple-choice response happens when the model is asked a question with the material for answering the question included. For example, the QA for ’Is hypertension a risk factor for cardiovascular disease, yes or no?’ is multiple-choice-orientated. In contrast, an open question without potential answers provided gives rise to a free-text response, an example QA being ’What are the common symptoms of influenza?’.
Task-specific FT Models BERT [70] 44.6 51.6 43.0 - -
RoBERTa [80, 81] 43.3 52.8 - - -
BioBERT [29, 90] 30.1 60.2 - - -
SciBERT [91] 29.5 57.4 - - -
ClinicalBERT [31] 29.1 49.1 - - -
BlueBERT [121, 227, 228] - 48.4 - - -
PubMedBERT [30] 38.1 55.8 - - -
General LLMs Med-PaLM-2 [11] 86.5 81.8 72.3 88.3 95.2
FLAN-PaLM (few-shot) [10] 67.6 79.0 57.6 - -
GPT-4 (zero-shot) [7] 78.9 75.2 69.5 86.0 93.0
GPT-4 (few-shot) [7] 86.1 80.4 73.7 86.4 93.8
GPT-3.5 (zero-shot) [6] 50.8 71.6 50.1 69.8 70.2
GPT-3.5 (few-shot) [6] 60.2 78.2 62.7 68.7 69.8
Clinical Camel-13B (zero-shot) [18] 34.4 72.9 39.1 54.0 51.8
Clinical Camel-13B (5-shot) [18] 45.2 74.8 44.8 60.4 53.3
Clinical Camel-70B (zero-shot) [18] 53.4 74.3 47.0 69.8 71.3
Clinical Camel-70B (5-shot) [18] 60.7 77.9 54.2 72.8 75.0
Galactica [85] 44.4 77.6 52.9 - -
BioMedLM [95] 50.3 74.4 - - -
BioGPT [94] - 81.0 - - -
PMC-LLaMA [13] 56.4 77.9 56.0 - -
Human (expert) [123, 229] 87.0 78.0 90.0 - -
@@
The commonly used biomedical QA datasets are MedQA (USMLE) [14], PubMedQA [133], and MedMCQA [131]. Such datasets are relatively small in size compared to general datasets, thus the model performance can be readily improved by pre-training on general data and then finetuning on the biomedical data [230, 231]. Pergola et al. [232] proposed a biomedical-specific masking method: instead of masking tokens randomly, the model identifies biomedical-related tokens and masks them to focus on in-domain learning. Wang et al. [233] defined the self-questioning prompting (SQP) and utilize it on the BioASQ dataset. To extract useful information for specific tasks, SQP lets the GPT model ask questions about the given text and then answer these questions. However, hallucination remains a threat to the quality of outputs in QA. One solution is to use biomedical search systems such as Almanac [234]. Another approach is to use extra datasets as an augmentation for the QA task. There are many chatbots based on LLM QA, such as Clinical Camel [18], DoctorGLM [32], ChatDoctor [15], HuaTuo [17], HuaTuoGPT [33], and MedAlpaca [16]. Except for a few [235], most chatbots are a black box the the consumers, and thus providing insight into them is a potential direction for further studies.
@@
Existing works use accuracy as the metric for model performance in QA. We compare the performance of task-specific (fine-tuned) BERT variations [30, 236, 14, 237], Med-PaLM-2 [11, 18], FLAN-PaLM [10], GPT-4 [238, 18], GPT-3.5 [116, 18], Clinical Camel [18], Galactica [85], BioMedLM [95], BioGPT [94], and PMC-LLaMA [13] on the datasets described in Table 2. The citations in this paragraph correspond to the sources that provide data for assessing the model performance. Table 3 details their performances on five widely-used datasets, indicating that GPT-series models perform much better than BERT-series models. For GPT-series models, the observation that the few-shot setting outperforms the zero-shot setting still holds. Med-PaLM-2 has the highest accuracy among all models.  These results indicate that it is feasible to fine-tune general LLMs on medical data and improve their performance.
@@
Types Models NCBI Disease BC5CDR Disease BC5CDR Drug/Chem. JNLPBA
Task-specific FT Models BERT [70] 85.63 82.41 91.16 74.94
BioBERT [29, 90] 89.36 86.56 93.44 77.59
SciBERT [91] 88.57 90.01 90.01 77.28
PubMedBERT [30] 87.82 85.62 93.33 79.10
BlueBERT [121, 227, 228] 88.04 83.69 91.19 77.71
General LLMs ClinicalBERT [31] 86.32 83.04 90.80 78.07
GPT-3 [6] 51.40 43.60 73.00 -
GPT-3.5 (zero-shot) [6] 24.05 - 29.25 -
GPT-3.5 (one-shot) [6] 12.73 - 18.03 -
GPT-4 (zero-shot) [7] 56.73 - 74.43 -
GPT-4 (one-shot) [7] 48.37 - 82.07 -
ChatGPT [239] 50.49 51.77 60.30 41.25
BARD (zero-shot) [87] 96.00 - 97.70 -
BARD (5-shot) [87] 95.60 - 98.30 -
@@ A.2 Entity Extraction
Entity extraction, or named entity recognition (NER) [240] aims to identify named entities mentioned in unstructured text, and classify them into predefined categories such as the names of persons, organizations, locations, medical codes, time expressions, quantities, monetary values, percentages, etc. Existing task-specific fine-tuned models leverage large-scale pre-trained language representations, and then fine-tune them on NER datasets to achieve state-of-the-art performances. The self-attention mechanisms [69] are used for efficiently capturing complex patterns and dependencies.
@@
Table 2 shows commonly used biomedical NER datasets, such as NCBI Disease [143], JNLPBA [144], and BC5CDR [146]. To perform the NER task, for each input token, the existing models output a dense representation, which not only embeds the tokens but also includes its relation with other tokens in the text. Therefore, with the dense representations of input tokens extracted as vectors, additional layers are applied to the last Transformer layer to fit the downstream entity extraction task. The widely-used additional layers are softmax, BiLSTM [241], CRF [242, 243], and their combinations [244, 242, 245, 243, 246, 247]. Gu et al. [248] proposed a distillation method that uses few-shot GPT-3.5 to extract the correct entities and create the training set for the student model. This method shows that GPT models can be used to label the dataset first to achieve unsupervised learning. Wang et al. [233] defined a new prompting method self-questioning prompting (SQP) for NER. Wang et al. [233] further evaluated the performances of BARD, GPT-3.5, and GPT-4 with SQP and achieved some state-of-the-art results.
@@
The entity-level F1 score is widely used to evaluate the models’ performance. Recent studies on the model performance of LLMs on biomedical NER tasks include GPT-3 [249, 250, 251], GPT-3.5 [62], GPT-4 [62], and ChatGPT [61]. The citations in this paragraph correspond to the sources that provide data for assessing the model performance.
We summarize the performances of existing LLMs on the NER task in Table 4. It clearly shows that both the encoder-only BERT-series models (e.g. BioBERT [29], SciBERT [91], PubMedBERT [30], BlueBERT [121], and ClinicalBERT [31]) significantly outperform most decoder-only LLM GPT-series models (e.g., GPT-3 [6] and GPT-4 [7]), under both the zero-shot and few-shot settings. Recent work has investigated the performance of GPT-3 [249, 250, 251], GPT-3.5 [62], and GPT-4 [62] on biomedical NER tasks. These GPT models still need further development to outperform traditional task-specific fine-tuned models [251].
@@ A.3 Relation Extraction
For relation extraction, Table 2 shows representative biomedical datasets such as BC5CDR, BioRED, and DDI. In practice, most datasets indicate the potential entity for the relation instead of asking the model to find it, so RE is usually considered as a classification task. One common technique for task-specific fine-tuned models is to use the classification token (e.g., [CLS] token in Transformer [69]), which learns the information of all input tokens, and is further used to classify the relation. Some studies use only the classification token and apply softmax in the last layer [252, 253, 254, 255]. Others use not only the classification token, but also all other token representations. It has been shown that using all token representations outperforms using classification token only [256]. Specifically, Wang et al. [233] defined the self-questioning prompting (SQP) as discussed in Sec A.1 and utilize it in the DDI RE task.
@@
Existing studies on RE model performance typically use the F1-score, since the model performs a straightforward classification task. Table 5 compares the performance of BioBERT [29, 257], SciBERT [91, 257], PubMedBERT [30, 257], BioGPT [94], GPT-3 [249, 250, 251], GPT-3.5 [62], GPT-4 [62], and ChatGPT [61] on the datasets introduced above. It clearly shows that at present task-specific fine-tuned models outperform GPT models. For GPT models, the F1-score of the best GPT model (GPT-4) is below 70%, while that of all the task-specific fine-tuned models based on BERT is approximately 80%. There is no appreciable difference in F1-score between zero-shot and few-shot [233].
@@
Types Models BC5CDR ChemProt DDI GAD
Task-specific FT Models BioBERT [29, 90] - 76.46 80.88 82.36
SciBERT [91] 79.00 83.64 84.08 81.34
PubMedBERT [30] - - 82.36 83.96
BioGPT [94] 46.17 - 40.76 -
General LLMs GPT-3 [6] - 25.90 16.10 66.00
GPT-3.5 (zero-shot) [6] - 57.43 33.49 -
GPT-3.5 (one-shot) [6] - 61.91 34.40 -
GPT-4 (zero-shot) [7] - 66.18 63.25 -
GPT-4 (one-shot) [7] - 65.43 65.58 -
ChatGPT [239] - 34.16 51.62 52.43
BARD (zero-shot) [87] - - 56.60 -
BARD (5-shot) [87] - - 77.20 -
@@
Task-specific FT Models PubMedBERT [30] 82.32 - - -
BioGPT [94] 85.12 - 82.30 -
General LLMs GPT-4 (5-shot) [7] - 99.00 67.40 -
ChatGPT (0-shot) [239] - 92.90 - 39.93
ChatGPT (2-shot) [239] - 92.90 - 47.05
ChatGPT (5-shot) [239] - 92.90 60.00 45.39
ChatGraph [258] - - - 60.79
@@
Task-specific FT Models BERT [259] 76.11
ALBERT [259] 77.84
Roberta [259] 80.14
BioBERT [259] 81.83
ClinicalBERT [259] 80.66
BlueBERT [259] 83.92
SciBERT [259] 79.43
BlueBERT+AG [259] 84.34
General LLMs GPT-3.5 [6] 82.21
GPT-3.5-Distillation [248] 80.24
GPT-4 [7] 85.69
BARD (zero-shot) [87] 76.00
BARD (5-shot) [87] 76.00

### v5

@@ Acknowledgements
Competing Interests
The authors declare no competing interests.
@@ Appendix A Appendix: Discriminative Tasks
Types Tasks Datasets Data Scale Data Source
Discriminative Tasks (Sec. A) Question Answering (Sec. A.1) MedQA (USMLE) [medqa] 61,097 multiple-choice QA pairs 51 medical textbooks
MedMCQA [Medmcqa] 194k multiple-choice QA pairs AIIMS and NEET PG entrance exam
MMLU [hendrycks2020measuring] 300 multiple-choice QA pairs Clinical Knowledge
PubMedQA [PubmedQA] 273,500 question-context-answer triples PubMed [PubMed, sen2008collective]
BioASQ-QA [krithara2023bioasq] 4,721 QA pairs BioASQ [tsatsaronis2015overview]
EMRQA [pampari2018emrqa] 400k+ QA pairs i2b2
CliCR [vsuster2018clicr] 105k QA pairs 12k clinical case reports
COVID-QA [moller2020covid] 2,019 QA pairs 147 articles (CORD-19 [wang2020cord])
MASH-QA [zhu2020question] 34,808 QA pairs WebMD website
Entity Extraction (Sec. A.2) NCBI Disease [dougan2014ncbi] 6,892 entity mentions 793 PubMed abstracts
BC5CDR [li2016biocreative] 28,785 entity mentions 1,500 PubMed articles
BC4CHEMD [krallinger2015chemdner] 84,355 entity mentions 10,000 PubMed abstracts
BioRED [luo2022biored] 20,419 entity mentions 600 PubMed abstracts
CMeEE [zhang2021cblue] 24,850 entity mentions 938 files
ADE [gurulingappa2012development] 21,160 entity mentions 2,972 documents
2012 i2b2 [sun2013evaluating] 30,690 entity mentions 310 discharge summaries
2014 i2b2/UTHealth [stubbs2015annotating] 57,744 entity mentions 1,304 longitudinal medical records
2018 n2c2 [henry20202018] 83,869 entity mentions 505 discharge summaries[johnson2016mimic]
Cadec [karimi2015cadec] 9,111 entity mentions 1,253 posts
Relation Extraction (Sec. A.3) BC5CDR [li2016biocreative] 3,116 relations 1,500 PubMed articles
BioRED [luo2022biored] 6,503 relations 600 PubMed abstracts
ADE [gurulingappa2012development] 13,806 relations 2,972 documents
2018 n2c2 [henry20202018] 59,810 relations 505 discharge summaries
2010 i2b2/VA [uzuner20112010] 1,748 reports clinical reports
DDI [segura2013semeval] 5,021 relations 1,017 biomedical texts
EU-ADR [van2012eu] 2,436 relations 100 abstracts
Text Classification (Sec. A.4) ADE [gurulingappa2012development] 20,967 sentences 2,972 documents
2014 i2b2/UTHealth [stubbs2015annotating] 1,304 records longitudinal medical records
HoC [baker2016automatic] 5,310 sentences 1,499 PubMed abstracts
OHSUMED [hersh1994ohsumed] 50,216 abstracts 50,216 medical abstracts
WNUT-2020 Task 2 [nguyen2020wnut] 10,000 tweets 10,000 tweets
Medical Abstracts [schopf2022evaluating] 14,438 abstracts 28,880 medical abstracts
MIMIC-III [johnson2016mimic] 52,724 discharge summaries 53,423 hospital admissions
Natural Language MedNLI [romanov2018lessons] 14,049 sentence pairs MIMIC-III [johnson2016mimic]
Inference (Sec. A.5) BioNLI [bastan2022bionli] 40,459 sentence pairs PubMed abstracts
Semantic Textual Similarity (Sec. A.6) MedSTS [wang2020medsts] 174,629 sentence pairs 3M clinical notes
2019 n2c2/OHNLP [wang20202019] 2,054 sentence pairs 113k clinical notes and MedSTS [wang2020medsts]
BIOSSES [souganciouglu2017biosses] 100 sentence pairs Text Analysis Conference (TAC)
Information Retrieval (Sec. A.7) TREC-COVID [voorhees2021trec] 8,691 3-level query-document relations CORD-19 [wang2020cord]
NFCorpus [boteva2016full] 712k 3-level query-document relations NutritionFacts.org (NF) website
BioASQ (BEIR) [thakur2021beir] 32,916 binary query-document relations BioASQ [tsatsaronis2015overview]
Generative Tasks (Sec. B) Text Summarization /Generation (Sec. B.1/Sec. B.3) MIMIC-CXR [johnson2019mimic] 128k Findings-Impression pairs radiology reports
PubMed [PubMed, sen2008collective] 36M+ citations and abstracts biomedical literature
PMC [PMC] 9,407,149 papers biomedical and life sciences literature
CORD-19 [wang2020cord] 140k+ papers COVID-19 literature
MentSum [sotudeh2022mentsum] 24,119 post-TLDR summary pairs 24,119 Reddit posts
MeQSum [abacha2019summarization] 1,000 question-summary pairs U.S. National Library of Medicine
Text Simplification MultiCochrane [joseph2023multilingual] 7,755 abstract-summary pairs Cochrane Library systematic reviews
(Sec. B.2) AutoMeTS [van2020automets] 3,300 sentence pairs English Wikipedia [kauchak2013improving]
Task-specific FT Models BERT [devlin2018bert] 44.6 51.6 43.0 - -
RoBERTa [yang2020measurement, van2020automets] 43.3 52.8 - - -
BioBERT [lee2020biobert, raj2019saama] 30.1 60.2 - - -
SciBERT [beltagy2019scibert] 29.5 57.4 - - -
ClinicalBERT [alsentzer2019publicly] 29.1 49.1 - - -
BlueBERT [peng2019transfer, mutinda2020detecting, mahajan2020identification] - 48.4 - - -
PubMedBERT [gu2021domain] 38.1 55.8 - - -
General LLMs Med-PaLM-2 [medpalm2] 86.5 81.8 72.3 88.3 95.2
FLAN-PaLM (few-shot) [medpalm] 67.6 79.0 57.6 - -
GPT-4 (zero-shot) [gpt-4] 78.9 75.2 69.5 86.0 93.0
GPT-4 (few-shot) [gpt-4] 86.1 80.4 73.7 86.4 93.8
GPT-3.5 (zero-shot) [GPT-3] 50.8 71.6 50.1 69.8 70.2
GPT-3.5 (few-shot) [GPT-3] 60.2 78.2 62.7 68.7 69.8
Clinical Camel-13B (zero-shot) [toma2023clinical] 34.4 72.9 39.1 54.0 51.8
Clinical Camel-13B (few-shot) [toma2023clinical] 45.2 74.8 44.8 60.4 53.3
Clinical Camel-70B (zero-shot) [toma2023clinical] 53.4 74.3 47.0 69.8 71.3
Clinical Camel-70B (few-shot) [toma2023clinical] 60.7 77.9 54.2 72.8 75.0
Galactica [taylor2022galactica] 44.4 77.6 52.9 - -
BioMedLM [venigalla2022biomedlm] 50.3 74.4 - - -
BioGPT [luo2022biogpt] - 81.0 - - -
PMC-LLaMA [wu2023pmc] 56.4 77.9 56.0 - -
Human (expert) [he2023survey, diao2023lmflow] 87.0 78.0 90.0 - -
@@ A.1 Question Answering
Question Answering (QA) [jin2022biomedical] aims to give answers to the given queries, by generating multiple-choice or free-text responses. A multiple-choice response happens when the model is asked a question with the material for answering the question included. For example, the QA for ’Is hypertension a risk factor for cardiovascular disease, yes or no?’ is multiple-choice-orientated. In contrast, an open question without potential answers provided gives rise to a free-text response, an example QA being ’What are the common symptoms of influenza?’.
@@
The commonly used biomedical QA datasets are MedQA (USMLE) [medqa], PubMedQA [PubmedQA], and MedMCQA [Medmcqa]. Such datasets are relatively small in size compared to general datasets, thus the model performance can be readily improved by pre-training on general data and then finetuning on the biomedical data [soni2020evaluation, yoon2019pre]. Pergola et al. [pergola2021boosting] proposed a biomedical-specific masking method: instead of masking tokens randomly, the model identifies biomedical-related tokens and masks them to focus on in-domain learning. Wang et al. [wang2023large] defined the self-questioning prompting (SQP) and utilize it on the BioASQ dataset. To extract useful information for specific tasks, SQP lets the GPT model ask questions about the given text and then answer these questions. However, hallucination remains a threat to the quality of outputs in QA. One solution is to use biomedical search systems such as Almanac [hiesinger2023almanac]. Another approach is to use extra datasets as an augmentation for the QA task. There are many chatbots based on LLM QA, such as Clinical Camel [toma2023clinical], DoctorGLM [xiong2023doctorglm], ChatDoctor [li2023chatdoctor], HuaTuo [wang2023huatuo], HuaTuoGPT [zhang2023huatuogpt], and MedAlpaca [han2023medalpaca]. Except for a few [li2023huatuo], most chatbots are a black box for consumers, and thus providing insight into them is a potential direction for further studies.
@@
Existing works use accuracy as the metric for model performance in QA. We compare the performance of task-specific (fine-tuned) BERT variations [gu2021domain, liu2023kbmqa, medqa, alghanmi2023interpreting], Med-PaLM-2 [medpalm2, toma2023clinical], FLAN-PaLM [medpalm], GPT-4 [nori2023capabilities, toma2023clinical], GPT-3.5 [lievin2022can, toma2023clinical], Clinical Camel [toma2023clinical], Galactica [taylor2022galactica], BioMedLM [venigalla2022biomedlm], BioGPT [luo2022biogpt], and PMC-LLaMA [wu2023pmc] on the datasets described in Table 4. The citations in this paragraph correspond to the sources that provide data for assessing the model performance. Table 5 details their performances on five widely-used datasets, indicating that GPT-series models perform much better than BERT-series models. For GPT-series models, the observation that the few-shot setting outperforms the zero-shot setting still holds. Med-PaLM-2 has the highest accuracy among all models.  These results indicate that it is feasible to fine-tune general LLMs on medical data and improve their performance.
@@
Types Models NCBI Disease BC5CDR Disease BC5CDR Drug/Chem.
Task-specific FT Models BERT [devlin2018bert] 85.63 82.41 91.16
BioBERT [lee2020biobert, raj2019saama] 89.36 86.56 93.44
SciBERT [beltagy2019scibert] 88.57 90.01 90.01
PubMedBERT [gu2021domain] 89.86 85.62 93.50
BioLinkBERT [yasunaga2022linkbert] 88.76 86.39 94.04
BlueBERT [peng2019transfer, mutinda2020detecting, mahajan2020identification] 88.04 83.69 91.19
General LLMs ClinicalBERT [alsentzer2019publicly] 86.32 83.04 90.80
GPT-3 [GPT-3] 51.40 43.60 73.00
GPT-3.5 (zero-shot) [GPT-3] 38.02 32.26 41.25
GPT-3.5 (few-shot) [GPT-3] 42.74 - 72.10
GPT-4 (zero-shot) [gpt-4] 57.85 - 81.90
GPT-4 (few-shot) [gpt-4] 58.39 - 82.43
ChatGPT [openai2023chatgpt] 50.49 51.77 60.30
BARD (zero-shot) [manyika2023overview] 96.00 - 97.70
BARD (few-shot) [manyika2023overview] 95.60 - 98.30
@@ A.2 Entity Extraction
Entity extraction, or named entity recognition (NER) [li2020survey] aims to identify named entities mentioned in unstructured text, and classify them into predefined categories such as the names of persons, organizations, locations, medical codes, time expressions, quantities, monetary values, percentages, etc. Existing task-specific fine-tuned models leverage large-scale pre-trained language representations, and then fine-tune them on NER datasets to achieve state-of-the-art performances. The self-attention mechanisms [transformer] are used for efficiently capturing complex patterns and dependencies.
@@
Table 4 shows commonly used biomedical NER datasets, such as NCBI Disease [dougan2014ncbi] and BC5CDR [li2016biocreative]. To perform the NER task, for each input token, the existing models output a dense representation, which not only embeds the tokens but also includes its relation with other tokens in the text. Therefore, with the dense representations of input tokens extracted as vectors, additional layers are applied to the last Transformer layer to fit the downstream entity extraction task. The widely-used additional layers are softmax, BiLSTM [Hochreiter1997LSTM], CRF [sun2021biomedical, yu2019biobert], and their combinations [portelli2021bert, sun2021biomedical, fraser2019extracting, yu2019biobert, chen2020using, kang2021umls]. Gu et al. [gu2023distilling] proposed a distillation method that uses few-shot GPT-3.5 to extract the correct entities and create the training set for the student model. This method shows that GPT models can be used to label the dataset first to achieve unsupervised learning. Wang et al. [wang2023large] defined a new prompting method self-questioning prompting (SQP) for NER. Wang et al. [wang2023large] further evaluated the performances of BARD, GPT-3.5, and GPT-4 with SQP and achieved some state-of-the-art results.
@@
The entity-level F1 score is widely used to evaluate the models’ performance. Recent studies on the model performance of LLMs on biomedical NER tasks include GPT-3 [agrawal2022large, caufield2023structured, gutierrez2022thinking], GPT-3.5 [chen2023large], GPT-4 [chen2023large], and ChatGPT [chen2023extensive]. The citations in this paragraph correspond to the sources that provide data for assessing the model performance.
We summarize the performances of existing LLMs on the NER task in Table 6. It clearly shows that both the encoder-only BERT-series models (e.g. BioBERT [lee2020biobert], SciBERT [beltagy2019scibert], PubMedBERT [gu2021domain], BlueBERT [peng2019transfer], and ClinicalBERT [alsentzer2019publicly]) significantly outperform most decoder-only LLM GPT-series models (e.g., GPT-3 [GPT-3] and GPT-4 [gpt-4]), under both the zero-shot and few-shot settings. Recent work has investigated the performance of GPT-3 [agrawal2022large, caufield2023structured, gutierrez2022thinking], GPT-3.5 [chen2023large], and GPT-4 [chen2023large] on biomedical NER tasks. These GPT models still need further development to outperform traditional task-specific fine-tuned models [gutierrez2022thinking].
@@ A.3 Relation Extraction
For relation extraction, Table 4 shows representative biomedical datasets such as BC5CDR, BioRED, and DDI. In practice, most datasets indicate the potential entity for the relation instead of asking the model to find it, so RE is usually considered as a classification task. One common technique for task-specific fine-tuned models is to use the classification token (e.g., [CLS] token in Transformer [transformer]), which learns the information of all input tokens, and is further used to classify the relation. Some studies use only the classification token and apply softmax in the last layer [wei2019relation, thillaisundaram2019biomedical, liu2020document]. Others use not only the classification token, but also all other token representations. It has been shown that using all token representations outperforms using classification token only [su2020investigation]. Specifically, Wang et al. [wang2023large] defined the self-questioning prompting (SQP) as discussed in Sec A.1 and utilize it in the DDI RE task.
@@
Existing studies on RE model performance typically use the F1-score, since the model performs a straightforward classification task. Table 7 compares the performance of BioBERT [lee2020biobert, wang2023pre], SciBERT [beltagy2019scibert, wang2023pre], PubMedBERT [gu2021domain, wang2023pre], BioGPT [luo2022biogpt], GPT-3 [agrawal2022large, caufield2023structured, gutierrez2022thinking], GPT-3.5 [chen2023large], GPT-4 [chen2023large], and ChatGPT [chen2023extensive] on the datasets introduced above. It clearly shows that at present task-specific fine-tuned models outperform GPT models. For GPT models, the F1-score of the best GPT model (GPT-4) is below 70%, while that of all the task-specific fine-tuned models based on BERT is approximately 80%. There is no appreciable difference in F1-score between zero-shot and few-shot [wang2023large].
@@
Types Models BC5CDR ChemProt DDI
Task-specific FT Models BioBERT [lee2020biobert, raj2019saama] - 76.46 80.88
SciBERT [beltagy2019scibert] 79.00 83.64 84.08
PubMedBERT [gu2021domain] - 77.24 82.36
BioGPT [luo2022biogpt] 46.17 - 40.76
BioLinkBERT [yasunaga2022linkbert] - 79.98 83.35
General LLMs GPT-3 [GPT-3] - 25.90 16.10
GPT-3.5 (zero-shot) [GPT-3] 43.29 57.43 33.49
GPT-3.5 (few-shot) [GPT-3] - 57.43 34.34
GPT-4 (zero-shot) [gpt-4] - 66.18 63.25
GPT-4 (few-shot) [gpt-4] - 66.82 65.58
ChatGPT [openai2023chatgpt] - 34.16 51.62
BARD (zero-shot) [manyika2023overview] - - 56.60
BARD (few-shot) [manyika2023overview] - - 77.20
@@
Task-specific FT Models PubMedBERT [gu2021domain] 82.34 - - -
BioGPT [luo2022biogpt] 85.12 - 82.30 -
General LLMs ChatGPT [openai2023chatgpt] 51.22 - - -
ChatGPT (zero-shot) [openai2023chatgpt] 65.72 92.90 - 39.93
ChatGPT (few-shot) [openai2023chatgpt] 69.32 92.90 60.00 47.05
GPT-4 (zero-shot) [gpt-4] 74.74 - - -
GPT-4 (few-shot) [gpt-4] 74.02 99.00 67.40 -
ChatGraph [shi2023chatgraph] - - - 60.79
@@
Task-specific FT Models BERT [wang2022paying] 76.11
ALBERT [wang2022paying] 77.84
Roberta [wang2022paying] 80.14
BioBERT [wang2022paying] 81.83
ClinicalBERT [wang2022paying] 80.66
BlueBERT [wang2022paying] 83.92
SciBERT [wang2022paying] 79.43
BlueBERT+AG [wang2022paying] 84.34
General LLMs GPT-3.5 [GPT-3] 82.21
GPT-3.5-Distillation [gu2023distilling] 80.24
GPT-4 [gpt-4] 85.69
BARD (zero-shot) [manyika2023overview] 76.00
BARD (few-shot) [manyika2023overview] 76.00